Find the complex zeros of each polynomial function. Write f in factored form.
f(x)= x^4-1

Answer by ewatrrr(24785)

 
Hi,
f(x)= x^4-1 Factor & Set to Zero
f(x)= (x^2+1)(x^2-1) = 0
x^2 = -1
x = i, -i Complex Roots
x^2 = 1
x = 1, -1